OpenStudy (anonymous):

Laplace Transform: 10(1-e^(-3t))u(t-1) ??

OpenStudy (anonymous):

Distribute the step function in first. Then for the term with the stepfunction and the exponential, use the formula L{ g(t)(u-a)} =e^-as L{g(t+a)}. So g(t+a) will be e^(-3+1)=e^(-2). So just take the laplace of that and multiple it by e^-s